# DualFluidNet: an Attention-based Dual-pipeline Network for Accurate and Generalizable Fluid-solid Coupled Simulation

### v1

@@ 3.2 Dual-pipeline Architectrue
Where $\phi$ represents the $CConv$ function(Equation 1). $S$ represents the Attention-based Particle Selector and its architecture is illustrated in the 3(b). With the concatenated features from the two pathways as input, the selector outputs a selection weight between 0 and 1, which is used to assess the respective significance of features from both pathways for fusion, aiming to achieve an optimal feature fusion. Unlike previous approaches, this kind of Attention-based fusion method allows our network to learn the optimal fusion between the two pathways autonomously, endowing our network with strong learning capabilities while ensuring compliance with the law of momentum conservation.
@@ 3.3 Type-aware Input Module (TaIM)
Figure 5: The schematic of the Type-aware Input Module in the Constraint-guided Path, with ASCC as the kernel. In contrast, the Main Path’s Type-aware Input Module replaces the ASCC kernel with CConv. The Particle Selector is the same as Figure 3(b).
We propose an innovative approach to facilitate better differentiation between fluid and solid particles in the input, promoting a better understanding of fluid-solid coupling. Consequently, this enables the network to more accurately and reliably compute their interactions during collisions. Figure 5 provides an example of the Type-aware Input Module architecture in the Constraint-guided Path, with ASCC as the kernel. In contrast, the Type-aware Input Module in the Main Path replaces the ASCC kernel with the CConv kernel. Unlike the Attention-based Feature Fusion Module between two pathways, the input features here represent two entirely different types, posing a greater challenge for our network to recognize their respective diverse properties. The quality of the initial fusion can significantly influence the final fusion weights. Since it is still a feature fusion problem, an intuitive approach is to employ an additional attention module for fusing these input features, which can be represented as:
@@
$f_{coup}^{2}=S(f_{coup}^{1})\otimes f_{fluid}+(1-S(f_{coup}^{1})\otimes f_{solid}.$ (6)
@@ 3.4 Training Procedure
$\mathcal{L}^{n+1}=\sum_{i=1}^{N}\varphi_{i}\left\|x_{i}^{t+1}-\hat{x}_{i}^{t+1}\right\|_{2}^{\gamma}.$ (7)
@@ 4 Experiments
Method Average pos err(mm) Average distance to Wasserstein Inference
n+1 n+2 closest point $d^{n}$(mm) distance time(ms)
Liquid(6k) DPI-Nets [13] 26.19 51.77 unstable - 305.55
KPConv [26] 1.65 4.54 unstable - 57.89
CConv [27] 0.56 1.51 29.50 0.26 16.47
DMCF [19] 0.65 1.89 32.89 0.21 94.86
Ours 0.43 1.16 28.32 0.17 48.01
DamBreak DPI-Nets [13] 12.73 25.38 22.07 - 202.56
KPConv [26] 2.49 7.05 unstable - 47.96
CConv [27] 0.57 1.38 20.63 0.21 12.01
DMCF [19] 0.64 1.83 24.56 0.20 70.49
Ours 0.42 1.02 20.90 0.18 35.57
Table 1: Quantitative Evaluation. We evaluated the networks according to the evaluation metrics described in 4.2, and also compared the inference time required per frame for each network. It’s worth noting that some methods become unstable after a few frames.
@@
Our experiments aimed to showcase the prediction accuracy, probability distribution divergence, and inference time of our approach compared to various baseline methods. What’s more, we show the great generalization and broad applicability of our network with adequate visual analysis.
4.1 Datasets
We conducted experiments on three distinct 3D fluid datasets. Two of these datasets were derived from the Liquid3D dataset presented by [27]. The first dataset, denoted as ’Liquid(6k),’ represents a simpler scenario with a fixed particle count of 6000. The second dataset, denoted as ’Liquid(complex),’ encompasses a variety of complex scenarios featuring different initial shapes of fluid particles and multiple complex boxes, with particle counts ranging from thousands to tens of thousands. The Liquid3D dataset was generated using DFSPH [2], which emphasizes high-fidelity simulations and operates at a 50 Hz time resolution. It comprises 200 training scenes and 20 test scenes. The third dataset, ’DamBreak,’ was from [13] and consists of 2000 training scenes and 300 testing scenes. These scenes simulate the behavior of a randomly positioned fluid block in a static box.
@@ 4.2 Evaluation metrics
To evaluate the prediction accuracy of each method, we calculate the average positional error by comparing the predicted positions to the ground truth. We start the assessment by using every $5^{th}$ frame for initialization and compute the deviations from the ground truth between two consecutive frames, denoted as $t+1$ and $t+2$. Furthermore, we measure long-term similarity by reporting the average distance from the ground-truth particles to the closest particle in the entire sequence. This distance for frame $n$ is computed as follows:
@@
Where $X^{n}$ represents the set of predicted particle positions for frame $n$, $\hat{x}_{i}^{n}$ corresponds to the ground-truth position for particle $i$, and $N$ denotes the total number of particles. We quantify the divergence between the prediction and ground truth probability distributions by calculating the Wasserstein Distance [23] between them. The Wasserstein Distance, also known as the Earth Mover’s Distance (EMD), is a measure of the transportation cost required to transform one probability distribution into another, which can be used to measure the similarity between two distributions.
@@ 4.3 Comparative Experiments
Method Average pos err(mm) Average distance to Wasserstein
t+1 t+2 closest point $d^{n}$(mm) distance
Main Path (no TaIM) 0.66 1.86 30.83 0.26
Main Path 0.58 1.61 30.67 0.21
Constraint-guided Path (no TaIM) 0.75 2.18 36.94 0.30
Constraint-guided Path 0.67 1.86 32.20 0.25
Ours 0.52 1.45 29.42 0.19
@@
Table 2: Ablation Study. Our ablation experiments were conducted on the ’Liquid (complex)’ dataset, testing the following configurations: a single Main Path without TaIM (Type-aware Input Module), a single Main Path with TaIM, a single Constraint-guided Path without TaIM, a single Constraint-guided Path with TaIM, and our final Dual-pipeline network.
Method Average pos err(mm) Average distance to Wasserstein
t+1 t+2 closest point $d^{n}$(mm) distance
CConv 0.66 1.86 30.83 0.26
CConv-TaIM 0.58 1.61 30.67 0.22
DMCF 0.73 2.08 47.89 0.21
DMCF-TaIM 0.69 2.03 46.58 0.21
Ours (no TaIM) 0.57 1.59 30.04 0.20
Ours 0.52 1.45 29.42 0.19
@@
Figure 7: Generalization to fluid initial shapes: the Stanford Bunny fluid free-falls into a box with a circular depression at the bottom, creating splashes. See the supplementary video.
Figure 8: Generalization to complex terrains and varying fluid particle scenarios: fluid emitted from a particle emitter flows from a higher elevation in a complex valley, eventually filling the depressions in a plain. See the supplementary video.
Quantitative Evaluation
Table 1 presents comparative experiments conducted on the Liquid(6k) and DamBreak datasets. We assess the average error between the ground-truth particle positions and the two predicted future frames in the test set. Additionally, we calculate the average distance from the ground truth to the prediction throughout the entire sequence. The Wasserstein Distance is employed to gauge the similarity between the predicted particle distribution and the ground truth distribution.
We compare our method to DPI-Nets [13], two kinds of continuous convolution network [26, 27] and DMCF [19]. It is worth noting that DMCF exhibits an increase in mean squared error (MSE) compared to CConv [27]. This is because the anti-symmetric convolution layer introduces a strong constraint, which makes the problem considerably more challenging for the network to solve. Merely appending the ASCC at the end made it more complex to fine-tune the network to its current state. As a trade-off, the network exhibits enhanced generalization. What’s more, the Wasserstein Distance to the ground truth has also shown a decrease, indicating a distribution that more closely approximates the physically accurate ground truth. Our network outperforms in both prediction accuracy and probability distribution divergence compared to previous models. This achievement is attributed to the design of our Dual-pipeline architecture, which attains equilibrium between global fluid control and adherence to physical law constraints.
Qualitative Evaluation
Figure 6 presents a visual comparison of our network with CConv and DMCF on the Liquid3D dataset. To better analyze the underlying physical principles during the dynamic simulation process, we selected five frames from the free-fall to collision with the container’s bottom for analysis. CConv starts to deform its structure from the second frame onwards, which significantly deviates from the expected physical laws of free fall. This is attributed to CConv’s lack of emphasis on the law of momentum conservation, allowing it to overfit gravity and negate it. This reduces the problem to a local per-particle problem. In contrast, DMCF exhibits an improved adherence to the laws of physics during the free fall phase, thanks to the inclusion of ASCC. However, DMCF’s approach of simply adding ASCC as the final output layer compromises the overall stability of the fluid. This instability becomes evident from the second frame, where the edges of the cube exhibit noticeable distortion. On the other hand, our network excels in simulating the fluid by providing a stable global context while maintaining a strong grasp of physical principles throughout the simulation process, bringing it closer to the ground truth.
@@ 4.5 Generalization
We validate the generalization capability of our network to unseen scenarios from two perspectives. Firstly, Figure 7 presents the generalization to irregular initial shapes (Stanford Bunny), which is completely unseen and full of challenges, as the initial shapes in the Liquid3D dataset are all regular. Secondly, we simulated a scenario where fluid, emitted from a particle emitter, flows from a high point in a complex valley terrain down to fill a depression in a plain, demonstrating the network’s generalization ability regarding scene geometry and the number of particles, as shown in Figure 8. The results demonstrate the strong generalization capabilities of our method in handling complex liquid shapes and terrains. This is attributed to our Dual-pipeline architecture, which effectively learns fluid dynamics while tightly adhering to physical laws.

### v2

@@ 3.2 Dual-pipeline Architectrue
Where $\phi$ represents the CConv function(Equation 1). $S$ represents the Particle Selector and its architecture is illustrated in Figure 4. With the concatenated features from the two pathways as input, the selector outputs a selection weight between 0 and 1, which is used to assess the respective significance of features from both pathways for fusion, aiming to achieve an optimal feature fusion. Unlike previous approaches, this kind of Attention-based fusion method allows our network to learn the optimal fusion between the two pathways autonomously, endowing our network with strong learning capabilities while ensuring compliance with the law of momentum conservation.
@@ 3.3 Type-aware Input Module (TaIM)
We propose an approach to facilitate better differentiation between fluid and solid particles in the input, promoting a better understanding of fluid-solid coupling. Consequently, this enables the network to more accurately and reliably compute their interactions during collisions. Figure 5 provides an example of the Type-aware Input Module architecture in the Constraint-guided Path, with ASCC as the kernel. In contrast, the Type-aware Input Module in the Main Path replaces the ASCC kernel with the CConv kernel. Unlike the Attention-based Feature Fusion Module between two pathways, the input features here represent two entirely different types, posing a greater challenge for our network to recognize their respective diverse properties. The quality of the initial fusion can significantly influence the final fusion weights. Since it is still a feature fusion problem, an intuitive approach is to employ an additional attention module for fusing these input features, which can be represented as:
@@
$f_{coup}^{2}=S(f_{coup}^{1})\otimes f_{fluid}+(1-S(f_{coup}^{1}))\otimes f_{solid}.$ (8)
@@ 3.4 Training Procedure
$\mathcal{L}^{t+1}=\sum_{i=1}^{N}\varphi_{i}\left\|x_{i}^{t+1}-\hat{x}_{i}^{t+1}\right\|_{2}^{\gamma}.$ (9)
@@ 4 Experiments
Our experiments aimed to showcase the particle prediction accuracy, probability distribution divergence, adherence of physical law, and inference time of our approach compared to various baseline methods. What’s more, we show the great generalization and broad applicability of our network with adequate visual analysis.
4.1 Experimental Setup
We utilize the PyTorch framework and employ the Adam optimizer for training, with the batch size set to 16. The initial learning rate is 0.002, and we halve the learning rate at steps 10000, 20000, …, 50000. In total, we train our networks for 50,000 iterations on an NVIDIA RTX 3090Ti. Most other parameters follow the settings in [39]. The particle radius is set to $h=0.025m$, and the spherical filters with a spatial resolution of [4,4,4] are used with a radius of $R=4.5h$.
Method Average pos error(mm) Average distance of Wasserstein Max density Frame inference
t+1 t+2 $n$ frames sequence $d^{n}$(mm) distance(mm) error($g/cm^{3}$) time(ms)
Liquid(6k) DFSPH [2] - - - - - $\geq 1,000$
DPI-Nets [23] 26.01 50.67 unstable - unstable 305.55
KPConv [37] 1.65 4.54 unstable - unstable 57.89
PCNN [43] 0.64 1.87 32.50 0.33 0.15 187.34
Regression-Forests [20] 0.68 1.97 38.47 0.29 0.13 12.98
CConv [39] 0.60 1.55 30.09 0.26 0.12 16.47
DMCF [29] 0.65 1.89 32.89 0.21 0.07 94.86
Ours 0.43 1.16 28.32 0.17 0.06 48.01
DamBreak DFSPH [2] - - - - - $\geq 1,000$
DPI-Nets [23] 12.70 24.38 22.57 - 0.16 202.56
KPConv [37] 2.49 7.05 unstable - unstable 47.96
PCNN [43] 0.75 1.93 25.43 0.31 0.14 187.34
Regression-Forests [20] 0.72 1.62 31.40 0.25 0.14 11.10
CConv [39] 0.57 1.38 20.63 0.22 0.11 12.01
DMCF [29] 0.64 1.83 24.56 0.20 0.07 70.49
Ours 0.42 1.02 20.09 0.18 0.06 35.57
Table 2: Quantitative Evaluation. We evaluated the networks according to the evaluation metrics described in Section 4.3 and also compared the inference time required per frame for each network. It’s worth noting that some methods become unstable after a few frames.
4.2 Datasets
To evaluate the fundamental fluid simulation capability, we conducted experiments on three basic 3D fluid datasets. Two of these datasets were derived from the Liquid3D dataset presented by [39]. The first dataset, denoted as ’Liquid(complex),’ encompasses a variety of complex scenarios featuring different initial shapes of fluid particles and multiple complex boxes, with particle counts ranging from thousands to tens of thousands. The second dataset, denoted as ’Liquid(6k),’ represents a simpler scenario with a fixed particle count of 6000. The Liquid3D dataset was generated using DFSPH [2], which emphasizes high-fidelity simulations and operates at a 50 Hz time resolution. It comprises 200 training scenes and 20 test scenes. The third dataset, ’DamBreak,’ was from [23] and consists of 2000 training scenes and 300 testing scenes. These scenes simulate the behavior of a randomly positioned fluid block in a static box. We validate and compare the fundamental fluid simulation capabilities of all methods across these three datasets.
Figure 6: An example scene of the Tank3D dataset
To further explore the network’s ability to handle more complicated scenes, we propose a new dataset, Tank3D. Tank3D is constructed following the same methodology as Liquid3D [39], generated using DFSPH [2]. It consists of 100 randomly generated training scenes and 10 test scenarios each featuring a tank with a complex shape comprising three baffles with numerous holes. In each scene, 4 to 16 fluid particles are randomly generated for free fall. Optional fluid shapes remain consistent with Liquid3D, described in [39], but undergo random scaling and rotation. An example scene is shown in Figure 6. This Tank3D dataset features significantly more complex, serving as an effective benchmark for evaluating the neural network’s adherence to physical laws, fluid-solid interaction capability, and generalization ability. In Section 4.4.3, we further investigate the network’s handling abilities in complex terrains using the Tank3D dataset.
@@ 4.2 Datasets
Figure 8: Evaluation of the irregular initial shapes: the Stanford Bunny fluid free-falls into a box with a circular depression at the bottom. The sequence includes an initial frame and four simulation frames. The four moments we selected include ’the first meet of surface’, ’half rabbit submerged’, ’splash hits maximum height’, and ’water calms at the final frame’.
@@ 4.3 Evaluation metrics
To evaluate the prediction accuracy of each method, we use the chamfer distance (CD) by calculating the average positional error between the predicted positions to the ground truth. We start the assessment by using every $5^{th}$ frame for initialization and compute the deviations from the ground truth between two consecutive frames, denoted as $t+1$ and $t+2$. Furthermore, we measure long-term similarity by reporting the average distance from the ground-truth particles to the closest particle in the prediction for the whole sequence of $n$ frames. This distance for frame $n$ is computed as follows:
@@
Where $X^{n}$ represents the set of predicted particle positions for frame $n$, $\hat{x}_{i}^{n}$ corresponds to the ground-truth position for particle $i$, and $N$ denotes the total number of particles.
We quantify the probability distribution divergence between the prediction and ground truth by calculating the Wasserstein distance [32]. The Wasserstein distance, also known as the Earth Mover’s Distance (EMD), is a measure of the transportation cost required to transform one probability distribution into another, which can be used to measure the similarity between two distributions.
Notably, the chamfer distance (Equation 12) effectively calculates the aggregate simulation error at the particle level, thereby it can serve as a metric of the overall control capability in fluid simulations. However, it might not adequately capture global shape discrepancies, especially when there are substantial deformations in the shapes (for instance, in fluid motion). We find that the Wasserstein distance, which measures the error in the overall probability distribution, serves as a more indicative metric of adherence to physical laws. To show our network can not only achieve precise global fluid control but also ensure adherence to physical law constraints, it must excel in both metrics compared to previous methods.
Maximum density is defined as the highest density value among all fluid particles. This metric correlates with the fluid’s incompressibility and provides an indication of a method’s stability. The max density error is calculated as follows:
$e=\left|1-\frac{\max_{i}\rho(x_{i})}{\max_{i}\rho(\hat{x}_{i})}\right|$ (13)
Here, $\rho$ represents the function for calculating the density of particle $i$, where $x_{i}$ denotes the predicted particle $i$ and $\hat{x}_{i}^{n}$ corresponds to the ground-truth particle $i$.
@@ 4.4 Comparative Experiments
4.4.1 Quantitative Evaluation
Table 2 presents comparative experiments conducted on the Liquid(6k) and DamBreak datasets. Average pos error and Average distance of $n$ frames sequence assess error at the particle level, with the former evaluating short-term fluid prediction (next two frames) and the latter assessing long-term prediction (n frames). Wasserstein distance evaluates prediction accuracy at the level of probability distribution, while Max density error calculates the error between the maximum density of predicted particles and the ground truth, which correlates with the fluid’s incompressibility and provides an indication of a method’s stability. Finally, Frame inference time measures the average speed of inference per frame.
DFSPH [2] is a traditional fluid simulation method, which exhibits accuracy close to ground truth. However, the computational demands of traditional methods are hundreds of times higher than data-driven methods. We compare the simulation results with both traditional machine learning [20] and deep learning approaches [23, 39, 29, 37, 43]. Traditional machine learning methods, exemplified by regression forests [20], boast rapid inference speeds. However, they exhibit significant shortcomings in stability and generalizability compared to other existing approaches, which will be further analyzed in Section 4.4.3. It is worth noting that, despite DMCF [29] being an improved version of CConv [39], it underperforms CConv in terms of the chamfer distance metric. This is because the anti-symmetric continuous convolution layer introduces a strong constraint, which makes the problem considerably more challenging for the network to solve. Merely appending the ASCC at the end in [29] made it more complex to learn and fit the global fluid control. As a trade-off, the network exhibits improved adherence to physical laws and enhanced generalization capabilities: the Wasserstein distance to the ground truth has also shown a decrease, indicating a distribution that more closely approximates the physically accurate ground truth. As for our network, while it may not be the fastest neural network, it outperforms in both prediction accuracy and probability distribution divergence compared to previous models, maintaining a small error of max density. A smaller Max density error signifies enhanced fluid incompressibility and stability. These achievements are attributed to the design of our Dual-pipeline architecture, which attains equilibrium between global fluid control and adherence to physical law constraints.
4.4.2 Qualitative Evaluation
Figure 7 presents a visual comparison of our network with the most stable methods CConv and DMCF on the Liquid3D(complex) dataset. To better analyze the underlying physical principles during the dynamic simulation process, we selected five frames from the free-fall to collision with the container’s bottom for analysis. The visual contrast may not be immediately striking at first glance, but the details inherently reveal the level of commitment to following physical laws. In detail, CConv starts to deform its structure from the second frame onwards, which significantly deviates from the expected physical laws of free fall. This is attributed to CConv’s lack of emphasis on the law of momentum conservation, allowing it to overfit gravity and negate it. This reduces the problem to a local per-particle problem. In contrast, DMCF exhibits an improved adherence to the laws of physics during the free fall phase, thanks to the inclusion of the antisymmetric design. However, DMCF’s approach of simply adding ASCC as the final output layer compromises the overall stability of the fluid. This instability becomes evident from the second frame, where the edges of the cube exhibit noticeable distortion. Our network demonstrates an accurate simulation by providing a stable global context while maintaining a strong grasp of physical principles throughout the simulation process, bringing it closer to the ground truth.
Figure 8 shows the results of the evaluation for irregular initial shapes (Stanford Bunny), which is completely unseen, as the initial shapes in the Liquid3D dataset are all regular. We selected five moments for comparison. In the face of complex initial fluid shapes, CConv, lacking a grasp of physical principles, resulted in the rabbit completely collapsing and deforming during its fall; On the other hand, DMCF more accurately reflected the physical properties of free fall but showed inadequacy in global fluid control. This is evident as some liquid particles splashed out of the container during the fall and collision, and the liquid reached a noticeably higher and more chaotic state than the ground truth at the peak moment. Our network, however, closely aligns with the ground truth throughout the free fall and collision, demonstrating its exceptional generalization capabilities in complex and unseen initial shapes.
Figure 9: Qualitative Evaluation on Tank3D Dataset. Comparative experiments in complex tank scenarios provide a deeper exploration of models’ capability to handle complex scenarios, emphasizing adherence to physical laws, fluid-solid interaction proficiency, and generalization capacity. The image in the top left corner represents the initial state at frame 0.
Figure 10: Evaluation on our method’s rotation invariance. Our approach maintains accurate predictions even for scenes that are randomly rotated and unseen before.
Method Average pos error(mm) Average distance of Wasserstein Max density
t+1 t+2 $n$ frames sequence $d^{n}$(mm) distance(mm) error($g/cm^{3}$)
Main Path (w/o TaIM) 0.62 1.80 30.65 0.26 0.12
Main Path 0.58 1.61 30.67 0.23 0.11
Constraint-guided Path (w/o TaIM) 0.75 2.18 36.94 0.30 0.15
Constraint-guided Path 0.67 1.86 32.20 0.25 0.12
Ours (w/o TaIM) 0.57 1.59 30.04 0.20 0.06
Ours 0.52 1.45 29.42 0.19 0.06
@@ 4.4.2 Qualitative Evaluation
Table 3: Ablation Study. Our ablation experiments were conducted on the ’Liquid (complex)’ dataset, testing the following configurations: a single Main Path with and without TaIM (Type-aware Input Module), a single Constraint-guided Path with and without TaIM, and our final Dual-pipeline network with and without TaIM.
Method Average pos error(mm) Average distance of Wasserstein Max density
t+1 t+2 $n$ frames sequence $d^{n}$(mm) distance(mm) error($g/cm^{3}$)
CConv 0.66 1.86 30.83 0.26 0.12
CConv(w/ TaIM) 0.59 1.69 30.70 0.23 0.11
DMCF 0.73 2.08 47.89 0.22 0.07
DMCF(w/ TaIM) 0.69 2.03 46.58 0.22 0.07
Ours (w/o TaIM) 0.57 1.59 30.04 0.20 0.06
Ours 0.52 1.45 29.42 0.19 0.06
@@
4.4.3 Capability of Handling Complex Scene
In the experimental results presented above, we employ the same experimental methods and datasets as existing approaches, investigating the fundamental capabilities of our proposed method and existing methods in fluid simulation tasks. To further explore the network’s ability to handle complicated scenes, we propose a dataset, Tank3D, inspired by the Liquid3D dataset. Although smaller in scale, it features significantly greater complexity.
Figure 11: Generalization to complex terrains and varying fluid particle scenarios: fluid emitted from a particle emitter flows from a higher elevation in a complex valley, eventually filling the depressions in a plain. The video is available in our GitHub repository.
We train the model for 50,000 iterations using the Tank3D dataset, followed by qualitative comparisons as shown in Figure 9. Regression forests [20] stand as a classic example of traditional machine learning methods applied to fluid simulation tasks. Like many machine learning techniques, they struggle to generalize well to scenarios not encountered in the training set. Another practical problem is that water surfaces often remain unsettled and fail to become perfectly flat, even by the final frame. CConv [39] lacks guidance from physical principles, leading to observable severe compression of the fluid in the first image, which significantly deviates from the incompressibility of fluids. Due to the kernel computation method of the SPH approach, intense mutual compression among fluid particles results in a drastic increase in the velocity calculations for each particle, causing the dispersion of particles seen in the second image. The performance of DMCF [29] is considerably more stable. However, it still demonstrates a deficiency in capturing the interactions between fluids and complex baffles within the tank, manifesting as the fluid disregarding the presence of baffles and flowing directly downwards. In contrast, our method yields simulation results that closely align with reality, both in terms of overall fluid control and adherence to physical laws.
Furthermore, Figure 10 demonstrates the rotation invariance of our method. Faced with scenes that have undergone random rotations and were not encountered in the training set, our method still produces accurate predictions. This indicates that our approach has genuinely learned the principles of fluid dynamics and adheres well to the physical laws of the real world. This is difficult to achieve for other methods, especially traditional machine learning approaches.
@@ 4.6 Generalization
We validate the generalization capability of our network to an unseen and complex scenario, where fluid, emitted from a particle emitter, flows from a high point in a complex valley terrain down to fill a depression in a plain, demonstrating the network’s generalization ability regarding scene geometry and the number of particles, as shown in Figure 11. The results demonstrate the strong generalization capabilities of our method in handling complex fluid shapes and terrains. This is attributed to our Dual-pipeline architecture, which effectively learns fluid dynamics while tightly adhering to physical laws.